# A non-degeneracy theorem for interacting electrons in one dimension

### v1

@@ Acknowledgements
The author thanks Chokri Manai, Timo Weidl, Robert van Leeuwen, and Markus Penz for fruitful discussions and helpful comments on this article. The author is also grateful to Rafael Antonio Lainez Reyes for pointing out the references on Sobolev spaces.
@@
Appendix A Quadratic forms and self-adjoint extensions
In this section we recall the basic theory of self-adjoint extensions via quadratic forms. The results presented here are standard and can be found in several classical references [RS75, RS78, AGHH88, Tes14]. We shall therefore omit the proofs, with an exception for part A.3 for which we have no reference.
A.1. Basic definitions
First, let us recall the definition of sesquilinear forms and self-adjoint operators.
Definition A.1 (Quadratic forms).
We say that $a:Q\times Q\rightarrow\mathbb{C}$ is a (densely defined) sesquilinear form in a Hilbert space $\mathcal{H}$ if $Q$ is a dense vector subspace of $\mathcal{H}$ and $a$ is antilinear in the first argument and linear in the second one. In this case, we call $Q$ the form domain of $a$. Moreover, we say that
@@ A.1. Basic definitions
    $a$ is symmetric if $a(\Psi,\Phi)=\overline{a(\Phi,\Psi)}$ for any $\Phi,\Psi\in Q$.
@@
    $a$ is semi-bounded on $\mathcal{H}$ if $a(\Psi,\Psi)\in\mathbb{R}$ and
    $a(\Psi,\Psi)\geq C\lVert\Psi\rVert_{\mathcal{H}}^{2},\quad\mbox{for any $\Psi\in Q$ and some $C\in\mathbb{R}$.}$
@@
    $a$ is closed in $\mathcal{H}$ if for any sequence $\{\Psi_{n}\}_{n\in\mathbb{N}}\subset Q$ satisfying
    $\lim_{n\rightarrow\infty}\Psi_{n}=\Psi\quad\mbox{in $\mathcal{H}$and}\quad\lim_{\max\{m,n\}\rightarrow\infty}a(\Psi_{n}-\Psi_{m},\Psi_{n}-\Psi_{m})=0,$
    it holds that $\Psi\in Q$ and $\lim_{n\rightarrow\infty}a(\Psi-\Psi_{n},\Psi-\Psi_{n})=0$.
Remark (Semibounded implies symmetry).
By definition any semibounded form satisfies $a(\Psi,\Psi)\in\mathbb{R}$ for $\Psi\in Q$. In particular, from the polarization identity, any semibounded form is automatically symmetric.
Remark (Closed and semibounded).
If $a$ is semibounded on $\mathcal{H}$ with lower bound $C$, then $a$ is closed if and only if $Q$ is a (complete) Hilbert space with respect to the norm
$\lVert\Psi\rVert_{a}=a(\Psi,\Psi)+(C+1)\lVert\Psi\rVert_{\mathcal{H}}^{2}.$ (A.1)
Definition A.2 (Self-adjoint).
We say that a linear operator $A:D(A)\subset\mathcal{H}\rightarrow\mathcal{H}$ is a semibounded self-adjoint operator if it satisfies the following conditions:
  1. (i)
    (Densely defined) $D(A)\subset\mathcal{H}$ is dense.
  2. (ii)
    (Symmetric) $\langle\Psi,A\Phi\rangle_{\mathcal{H}}=\langle A\Psi,\Phi\rangle_{\mathcal{H}}$, for any $\Psi,\Phi\in D(A)$.
  3. (iii)
    (Semibounded) there exists $C>0$ such that
    $\langle\Psi,A\Psi\rangle_{\mathcal{H}}\geq C\lVert\Psi\rVert_{\mathcal{H}}^{2},\quad\mbox{for any $\Psi\in D(A)$.}$
@@
    (Self-adjoint) $D(A^{\ast})=D(A)$, where
    $D(A^{\ast})\coloneqq\{\Psi\in\mathcal{H}\quad\mbox{such that thelinear map}\quad\Phi\in D(A)\mapsto\langle\Psi,A\Phi\rangle_{\mathcal{H}}\quad\mbox{is bounded in $\mathcal{H}$}\}.$
A.2. Quadratic forms and self-adjoint operators
The main result on the theory of self-adjoint extensions via quadratic forms is the following. For a proof, see [RS75].
Theorem A.3 (Equivalence between quadratic forms and self-adjoint operators).
Let $A$ be a positive self-adjoint operator on a Hilbert space $\mathcal{H}$, then the domain $D((A+1)^{\frac{1}{2}})\subset\mathcal{H}$ is dense and the sesquilinear form
$a_{A}(\Psi,\Phi)=\langle(A+1)^{\frac{1}{2}}\Psi,(A+1)^{\frac{1}{2}}\Phi\rangle-\langle\Psi,\Phi\rangle$
is a densely defined, symmetric, positive, and closed quadratic form on $\mathcal{H}$. Conversely, for any densely defined, positive, symmetric and closed quadratic form $a$, there exists a unique self-adjoint positive operator $A$ such that $a_{A}=a$.
Remark (Associated quadratic form).
We call $a_{A}$ the associated form of $A$.
Remark (The shifting trick).
The above construction works for any semibounded operator $H$ by first shifting $H$ by a large enough constant such that $A=H+C$ is positive, taking the associated form of $A$, and then shifting it back to $a_{H}=a_{A}-C\langle\cdot,\cdot\rangle_{\mathcal{H}}$. In particular, Theorem A.3 a one-to-one correspondence between semibounded self-adjoint operators and semibounded closed forms.
We now state the celebrated KLMN theorem [RS75, Theorem X.17]. For this, we recall the following definition.
Definition A.4 (Relatively bounded).
Let $a$ be a closed semibounded form on a Hilbert space $\mathcal{H}$ with form domain $Q$. Then we say that a symmetric sesquilinear form $b:Q\times Q\rightarrow\mathbb{C}$ is $a$-bounded if there exists $c,C>0$ such that
$|b(\Psi,\Psi)|\leq ca(\Psi,\Psi)+C\lVert\Psi\rVert_{\mathcal{H}}^{2},\quad\mbox{for any $\Psi\in Q$.}$ (A.2)
The infimum over all $c>0$ for which (A.2) holds is called the $a$-bound of $b$. Moreover, if this infimum is zero, then we say that $b$ is infinitesimally $a$-bounded.
Theorem A.5 (KLMN).
Let $a$ be a closed semibounded form on a Hilbert space $\mathcal{H}$ and suppose that $b$ is a relatively $a$-bounded with relative bound $<1$. Then the form $a+b$ is closed and semibounded, and therefore, there exists a unique self-adjoint operator associated to this form. Moreover, the norm in (A.1) induced by $a+b$ is equivalent to the norm induced by $a$.
Remark A.6 (Discrete spectrum and form domain).
A simple but useful consequence of the KLMN theorem is that, if the form domain of $a$ is compactly embedded in $\mathcal{H}$, then the self-adjoint operator associated to $a+b$ has purely discrete spectrum. In particular, this operator has an $\mathcal{H}$-orthonormal basis of eigenfunctions.
A.3. Pull-back
We now introduce the notion of the pull-back of a form via a linear transformation. We then collect some elementary properties of the pull-back map.
Definition A.7 (Pull-back).
Let $T:Q\rightarrow Q^{\prime}$ be a linear bounded operator between Hilbert spaces $Q$ and $Q^{\prime}$. Then for any quadratic form $a$ with form domain $Q^{\prime}$, we define the pull-back of $a$ via $T$ as the form
$T^{\#}a:Q\times Q\rightarrow\mathbb{C}\quad(T^{\#}a)(\Psi,\Phi)=a(T\Psi,T\Phi).$
It is immediate to verify that the pull-back preserves symmetry. The next lemma gives sufficient condition on $T$ for the pull-back to also preserve semiboundeness and closedness.
Lemma A.8 (Pull-back properties).
Let $T:\mathcal{H}\rightarrow\mathcal{H}^{\prime}$ be a bounded linear operator such that $T(Q^{\prime})\subset Q$ for some dense subspaces $Q\subset\mathcal{H}$ and $Q^{\prime}\subset\mathcal{H}^{\prime}$. Then the pull-back of any semibounded form in $\mathcal{H}^{\prime}$ with form domain $Q^{\prime}$ is a semibounded form in $\mathcal{H}$ with form domain $Q$. Moreover, if $T$ is bijective and $T(Q)=Q^{\prime}$, then the pull-back of any closed form with form domain $Q^{\prime}$ is a closed form with form domain $Q$.
@@ A.3. Pull-back
For the first part note that if $a$ is semibounded in $\mathcal{H}^{\prime}$ and $T$ bounded from $\mathcal{H}$ to $\mathcal{H}^{\prime}$, then
$(T^{\#}a)(\Psi,\Psi)=a(T\Psi,T\Psi)\geq C\lVert T\Psi\rVert_{\mathcal{H}^{\prime}}^{2}\geq\min\{0,C\}\lVert T\rVert_{\mathcal{H}\rightarrow\mathcal{H}^{\prime}}\lVert\Psi\rVert_{\mathcal{H}}^{2}.$
Thus $T^{\#}a$ is semibounded in $\mathcal{H}$. For the second part, let $\Psi_{n}\subset Q$ be a sequence such that $\Psi_{n}\rightarrow\Psi$ in $\mathcal{H}$ and
$\lVert\Psi_{n}-\Psi_{m}\rVert_{T^{\#}a}=(T^{\#}a)(\Psi_{n}-\Psi_{m},\Psi_{n}-\Psi_{m})\rightarrow 0\quad\mbox{as}\quad\max\{m,n\}\rightarrow\infty.$
Then, $T\Psi_{n}\rightarrow T\Psi$ in $\mathcal{H}^{\prime}$ and $a(T\Psi_{n}-T\Psi_{m},T\Psi_{n}-T\Psi_{m})\rightarrow 0$ as $m,n\rightarrow\infty$. As $a$ is closed, this implies that $T\Psi\in Q$ and $a(T\Psi-T\Psi_{n},T\Psi-T\Psi_{n})\rightarrow 0$. Since $T$ is bijective and $T(Q)=Q^{\prime}$, we have $\Psi=T^{-1}T\Psi\in Q^{\prime}$ and therefore $(T^{\#}a)$ is closed. ∎

### v2

@@ Acknowledgements
The author wishes to thank Chokri Manai for several inspiring discussions, which led to the ideas in the proof of Theorem 5.1. The author also thanks Timo Weidl, Robert van Leeuwen, and Markus Penz for fruitful discussions and helpful comments on this article. The author is also grateful to Rafael Antonio Lainez Reyes for pointing out the references on Sobolev spaces. In addition, the author is grateful to the anonymous referees for several suggestions to improve the presentation of the paper.
@@
Data availability
No datasets were generated or analysed during the current study.
Competing interests
The authors has no competing interests to declare that are relevant to the content of this article.
Appendix A Schrödinger operators with distributional potentials
In this section, we present the construction of self-adjoint realizations of $H_{N}(v,w)$ via their associated quadratic form. We also list some important examples of distributional potentials for which this construction holds. A detailed exposition on the theory quadratic forms and self-adjoint extensions can be found in several classical references [RS75, RS78, AGHH88, Tes14].
A.1. Self-adjoint realization
$|v(\rho_{\Psi,\Psi})+w(\rho^{(2)}_{\Psi,\Psi})|\leq\epsilon\lVert\Psi\rVert_{\mathrm{H}^{1}}^{2}+C_{\epsilon}\lVert\Psi\rVert_{\mathrm{L}^{2}}^{2},\quad\mbox{for any $\Psi\in\mathrm{H}^{1}(I_{N})$,}$ (A.1)
and any $\epsilon>0$ for some constant $C_{\epsilon}=C(v,w,\epsilon)>0$. Hence, from the celebrated KLMN theorem (see [RS75, Theorem X.17]), for any closed subspace $B\subset\mathrm{H}^{1/2}(\partial\Omega)$, we can construct a unique self-adjoint operator associated to the sesquilinear form
$\mathfrak{a}_{v,w}(\Psi,\Phi)=\int_{I_{N}}\overline{\nabla\Psi(x)}\boldsymbol{\cdot}\nabla\Phi(x)\mathrm{d}x+v(\rho_{\Psi,\Phi})+w(\rho^{(2)}_{\Psi,\Phi}),\quad\mbox{for $\Psi,\Phi\in\mathrm{H}^{1}_{B}(I_{N})\cap\mathcal{H}_{N}$.}$ (A.2)
Precisely, the following holds.
Lemma A.1 (Schrödinger operators with distributional potentials).
Let $N\in\mathbb{N}$, $v\in\mathcal{V}$, and $w\in\mathcal{W}$. Moreover, let $B\subset\mathrm{H}^{1/2}(\partial I_{N})$ be a closed subspace and $\mathrm{H}^{1}_{B}(I_{N})$ the trace-restricted Sobolev space introduced in Definition 3.8. Then the quadratic form in (A.2) is closed, symmetric and semibounded. In particular, there exists a unique self-adjoint operator $H_{N}^{B}(v,w)$ associated to this form. Moreover, $H_{N}^{B}(v,w)$ is semibounded and has purely discrete spectrum.
Remark A.2 (Discrete spectrum via compact form domain).
Recall that $H_{N}^{B}(v,w)$ has discrete specturm if and only if the form domain is compactly embedded in $\mathcal{H}_{N}$. Hence, the discrete spectrum statement in Lemma A.1 follows from the compact embedding $\mathrm{H}^{1}_{B}(I_{N})\subset\mathrm{H}^{1}(I_{N})\hookrightarrow\mathrm{L}^{2}(I_{N})$.
Remark A.3 (Notation).
Throughout the paper, we denote by $H_{N}(v,w;\Gamma)$ the self-adjoint realization of $\mathfrak{a}_{v,w}$ with form domain $\mathrm{H}^{1}(I_{N};\Gamma)\cap\mathcal{H}_{N}$ and by $H^{\alpha}_{N}(v;w;\Gamma)$ the realization with form domain $\mathrm{H}^{1}_{\alpha}(I_{N};\Gamma)\cap\mathcal{H}_{N}$.
A.2. From single-particle to many-particle boundary conditions
To this end, we start with the following lemma.
Lemma A.4 (Many-particle form domain).
Let $\alpha\in\mathbb{R}$ and $\mathrm{H}^{1}_{\alpha}(I)=\{f\in\mathrm{H}^{1}(I):f(0)=\alpha f(1)\}$. Then for any $N\in\mathbb{N}$, the span of Slater determinants
is dense in $\mathcal{Q}_{N}^{\alpha}=\mathrm{H}^{1}_{\alpha}(I_{N})\cap\mathcal{H}_{N}$.
Proof.
We first observe that the projection on the antisymmetric space, explicitly given in (5.3), is continuous with respect to the $\mathrm{H}^{1}$ topology and maps $N$-fold tensor products to $N$-fold Slater determinants. Therefore, it suffices to show that the space
$\alpha\partial_{x}\varphi_{j}(0)-\partial_{x}\varphi_{j}(1)=0.$
As a consequence of the previous lemma, we have the following relation between the single-particle operator $h_{\alpha}(v)\coloneqq H_{1}^{\alpha}(v,0)$ and the $N$-particle operator $H_{N}^{\alpha}(v,0)$ associated to the form
$\mathfrak{a}_{v}(\Psi,\Psi)=\int_{I_{N}}|\nabla\Psi(x)|^{2}\mathrm{d}x+v(\rho_{\Psi}),\quad\Psi\in\mathrm{H}^{1}_{\alpha}(I_{N})\cap\mathcal{H}_{N}.$
Lemma A.5 (From single-particle to non-interacting many-particle systems).
Let $v\in\mathcal{V}$, $\alpha\in\mathbb{C}$, and $N\in\mathbb{N}$. Let $\{\psi_{j}\}_{j\in\mathbb{N}}$ be an orthogonal basis of eigenfunctions [^2] Note that such a basis exists because the spectrum of $h_{\alpha}(v)$ is purely discrete, see Remark A.2. of $h_{\alpha}(v)=H_{1}^{\alpha}(v,0)$ with corresponding eigenvalues $\{\lambda_{j}\}_{j\in\mathbb{N}}$ ordered in non-decreasing order. Then the Slater determinants
$\{\Psi_{J}=\psi_{j_{1}}\wedge...\wedge\psi_{j_{N}}:J=\{j_{1},...,j_{N}\}\in\mathbb{N}^{N}\quad\mbox{such that}\quad 1\leq j_{1}<j_{2}...<j_{N}\}$ (A.3)
Proof.
From a long but straightforward calculation we find that
$\mathfrak{a}_{v}(\Psi_{J},\varphi_{1}\wedge...\wedge\varphi_{N})=\lambda_{J}\langle\Psi_{J},\varphi_{1}\wedge...\wedge\varphi_{N}\rangle_{\mathrm{L}^{2}},\quad\mbox{for any $\{\varphi_{j}\}_{j=1}^{N}\subset\mathrm{H}^{1}_{\alpha}(I)$,}$
where $\Psi_{J}=\psi_{j_{1}}\wedge...\wedge\psi_{j_{N}}$ and $\lambda_{J}=\sum_{k=1}^{N}\lambda_{j_{k}}$. Thus by Lemma A.4 and approximation we have
$\mathfrak{a}_{v}(\Psi_{J},\Phi)=\lambda_{J}\langle\Psi_{J},\Phi\rangle\quad\mbox{for any $\Phi\in\mathrm{H}^{1}_{\alpha}(I_{N})\cap\mathcal{H}_{N}$.}$
A.3. Examples of distributional potentials
To illustrate how large the class of admissible external and interaction potentials in Lemma A.1 is, let us present a few important examples.
  1. (1)
  2. (2)
    The $\delta$-interaction potential $w=\delta(x-y)$, defined via
    $w_{\delta}(\rho^{(2)}_{\Psi,\Phi})=\int_{I}\rho^{(2)}_{\Psi,\Phi}(x,x)\mathrm{d}x.$
  3. (3)
    belongs to $\mathrm{W}^{-1,q}(I_{2})$.
  4. (4)
    Lemma A.1 can also be extended to the case of $k^{th}$-body distributional potentials for $k\geq 3$. To be precise, for any $w_{k}\in\mathrm{W}^{-1,q}(I_{k})$ with $q>k$, one can use Lemma 3.13 and Young’s inequality to show that the form $(\Psi,\Phi)\mapsto w_{k}(\rho^{(k)}_{\Psi,\Phi})$, where
    $\rho^{(k)}_{\Psi,\Phi}(x_{1},..,x_{k})=\frac{N!}{(N-k)!k!}\int_{I_{N-k}}\overline{\Psi(x_{1},...,x_{N})}\Phi(x_{1},...,x_{N})\mathrm{d}x_{k+1}...\mathrm{d}x_{N},$
    is $\Delta$-bounded with relative bound 0. Therefore, any Schrödinger operator of the form
    is also included for $N\geq 6$.
Appendix B A prototypical example: the one-dimensional Laplacian
In this section we gather some well-known formulas for the eigenfunctions of the one-dimensional Laplacian under different boundary conditions. These examples serve to illustrate the results in Theorems 2.1, 2.3, and 2.6 and to justify Remark 2.4.
Throughout this section we denote respectively by $\Delta$, $\Delta_{0}$, and $\Delta_{\pm}$ the self-adjoint realizations of the one-dimensional Laplacian on $I$ with Neumann, Dirichlet, and periodic/anti-periodic boundary conditions. Precisely, these are the unique self-adjoint operators associated to the form
$\mathfrak{a}_{0}(\psi,\psi)=\lVert\nabla\psi\rVert_{\mathrm{L}^{2}}^{2},$
with form domains $\mathcal{Q}=\mathrm{H}^{1}(I)$, $\mathcal{Q}_{0}=\mathrm{H}^{1}_{0}(I)$, and $\mathcal{Q}_{\pm}=\mathrm{H}^{1}_{\pm 1}(I)$. The eigenvalues and eigenfunctions of these operators can be explicitly computed.
Proposition B.1 (Spectrum of one-dimensional Laplacian).
The following holds:
@@ Appendix B A prototypical example: the one-dimensional Laplacian
    (Neumann) We have
    $\sigma(-\Delta)=\{\pi^{2}(k-1)^{2}:k\in\mathbb{N}\}.$
    Moreover, each eigenvalue is simple and the associated normalized eigenfunction are
    $\varphi_{k}(x)=\sqrt{2}\cos(\pi(k-1)x),\quad k\in\mathbb{N}\quad\mbox{and}\quad\varphi_{1}(x)=1.$
@@
    (Dirichlet) We have
    $\sigma(-\Delta_{0})=\{\pi^{2}k^{2}:k\in\mathbb{N}\}.$
    Moreover, each eigenvalue is simple and the associated normalized eigenfunctions are
    $\varphi_{k}(x)=\sqrt{2}\sin(\pi kx).$
@@
    (Periodic) We have
    $\sigma(-\Delta_{+})=\{4\pi^{2}(k-1)^{2}:k\in\mathbb{N}\}.$
    Moreover, the lowest eigenvalue $\lambda_{1}=0$ is simple and the rest is double degenerate with the following normalized eigenfunctions
    $\varphi_{1}(x)=1,\quad\mbox{and}\quad\varphi_{k,1}(x)=\sqrt{2}\cos(2\pi(k-1)x),\quad\varphi_{k,2}(x)=\sqrt{2}\sin(2\pi(k-1)x).$
@@
    (Anti-periodic) We have
    $\sigma(-\Delta_{-})=\{\pi^{2}(2k-1)^{2}:k\in\mathbb{N}\}.$
    Moreover, every eigenvalue is double degenerate with normalized eigenfunctions
    $\varphi_{k,1}(x)=\sqrt{2}\cos(\pi(2k-1)x)\quad\mbox{and}\quad\varphi_{k,2}(x)=\sqrt{2}\sin(\pi(2k-1)x).$
@@
These results are classical and follow from straightforward calculations. ∎
One can now verify that Proposition B.1 is optimal in the following sense. By ordering the eigenvalues of the Laplacian in non-decreasing order and counting multiplicity $\{\lambda_{k}\}_{k\geq 1}$, we see that
$\lambda_{k}<\lambda_{k+1},$
holds if and only if $k$ is odd in the periodic case and $k$ is even in the anti-periodic case. Therefore, by considering the associated $N$-particle system of electrons, i.e., the operator $\Delta_{\pm}$ associated to $\lVert\nabla\Psi\rVert_{\mathrm{L}^{2}}^{2}$ with form domain $\mathcal{H}_{N}\cap\mathrm{H}^{1}_{\pm 1}(I_{N})$, we obtain explicit examples where the ground-state of the $N$-particle system is non-degenerate if and only if condition (2.5) holds.
In addition, note that by perturbing the Laplacian via external potentials, we can break the symmetry and make the eigenvalues of $\Delta_{\pm}$ simple. For instance, by adding $v=\alpha\delta_{x}$ for $\alpha>0$ small enough we can make the first eigenvalue of $\Delta_{-}$ or the second eigenvalue of $\Delta_{+}$ become simple. In particular, considering the $N$-particle version of such operators, we find systems where the ground-state is non-degenerate even though the condition in (2.5) is not fulfilled. This justifies Remark 2.4.